OpenStudy (anonymous):

5 SQUAREOOT OF NEGATIVE 100

OpenStudy (oaktree):

We start by evaluating the square root of negative one hundred. We do that like this:\[\sqrt{-100} = \sqrt{-1} \times \sqrt{100} = i \times 10 = 10i\]Then we multiply back the 5 on the oustide to get \[5 \times 10i = 50i\]Makes sense?